OpenStudy (anonymous):

convert 3.56 kg/m^3 to mg/cm^3

OpenStudy (koikkara):

\(\color{darkgreen}{~\huge\star~``Welcome~to~Open~Study~!!"~\star~}\) @bigbangchu

OpenStudy (koikkara):

1 cm^3 = 1 ml. [ (35.6 mg / ml) / (10^6 mg / kg) ] (10^3 mL / L) = 35.6 / 10^3 kg / L = 3.56 * 10^(- 2) kg / L. \(Hope~it~helps~!!\) Don't Forget, \(Appreciate~the~Best~Answer\) with a click on \(\color{purple}{"Best Response"}\) Button Too ! \(\color{darkgreen}{~''Nice~To~Meet~You~!!"}\)